A potentiometer wire of length 100 cm has  aresistance 10 ohm. It is connected in series with a tesistance and an accumulator of emf  20V and of negligible resistance. A source of emf 10 mV is balanced against a length of 40 cm of the potentiometer wire.What is the value of the external resistance. 

Grade:12

1 Answers

Ramesh V
14 years ago

Let ext. resistance be R ohma

let current be I

so 20 = I(10+R)

also I*4 = .01 V

so, I=2.5*10-3

on substitution, we get R = 7990 ohms

--

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult to understand - post it here and
we will get you the answer and detailed solution very quickly.We are all IITians and here to help you in your IIT JEE preparation. All the best.

Regards,
Naga Ramesh
IIT Kgp - 2005 batch